# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ 4.1. Implementation
We implement WW-FL based on the CrypTen MPC framework developed by Meta (Knott et al., 2021). CrypTen provides a PyTorch-style interface and implements MPC operations with GPU support. Specifically, CrypTen implements semi-honest arithmetic and Boolean secure two- and multi-party computation protocols that use a trusted third “helper” party to generate correlated randomness. CrypTen provides a “simulation” mode where the specified computation is performed on a single node in plaintext yet simulates all effects that computation in MPC would have on the results (e.g., due to limited fixed-point precision and truncation). We leverage this mode to efficiently evaluate WW-FL’s accuracy and later the impact of data-poisoning attacks. For our run-time and communication measurements, we benchmark the full MPC protocols. In all our experiments, the fixed-point precision in CrypTen is set to 22 decimal bits (the maximum value recommended by the developers).
@@
Setup. Plaintext FL and CrypTen-based WW-FL simulations are run on a single computing platform with two Intel Xeon Platinum 8168 CPUs, 1.5TB RAM, and 16 NVIDIA Tesla V100 GPUs. For realistic results in a distributed MPC deployment with two computational and one helper party, we use three Amazon AWS g3s.xlarge instances with 4 vCPUs, and 8GB GPU memory on an NVIDIA Tesla M60. These instances are located in the same AWS availability zone (due to high costs associated with routing traffic between different zones), yet we simulate intra- and inter-continental network connections by setting the bandwidth and latency to 1Gbps/100Mbps and 20ms/100ms, respectively.
Tasks. Following prior work such as (Ben-Itzhak et al., 2022), we evaluate WW-FL on two standard image classification tasks: recognizing (a) hand-written digits using LeNet trained on MNIST (LeCun et al., 1998), and (b) objects in one of 10 classes using ResNet9 trained on CIFAR10 (Krizhevsky, 2009). We simulate 1000 clients overall from which 100 are randomly sampled during each training iteration. In the FL setting, all sampled clients provide their computed update directly to the aggregation server. In the WW-FL setting, each of the 10 MPC clusters has 100 associated clients from which each samples 10 clients at random. Each client has 200 data points assigned to it at random with duplicates allowed between clients.
For plain FL, we use batch size 8 and learning rate 0.005, and train locally for 5 epochs before central aggregation. To allow for a fair comparison between FL and WW-FL, we use a correspondingly scaled batch size of 80 and a learning rate of 0.05 as in (Goyal et al., 2017). Each client and MPC cluster performs 5 local training epochs with either a batch size of 8 or 80, respectively, following the recommendation in (Goyal et al., 2017). Tab. 6 summarises all the (hyper) parameters used during the training.
Table 6. Training parameters used in FL and WW-FL.
Parameter FL WW-FL
@@ 4.2. Results
    What is the accuracy difference between FL (McMahan et al., 2017) and WW-FL (in plaintext)?
@@
    What is the impact on accuracy for WW-FL when moving from plaintext to (simulated) MPC?
@@
Q1 – FL vs WW-FL. In Fig. 3, we compare the validation accuracy of FL and WW-FL for image classification tasks for 500 rounds. Here, WW-FL converges significantly faster than regular FL, e.g., after 500 rounds of training ResNet9 on CIFAR10, WW-FL reaches 85.68% validation accuracy, whereas regular FL only reaches 65.95%. Similarly, for MNIST, WW-FL achieves 98.95% accuracy compared to 98.72% for regular FL. We attribute this to WW-FL pooling training data at the cluster level and thus being able to exploit the known benefits of batching (Goyal et al., 2017; Bottou et al., 2018). Further plots for up to 2000 epochs are given in Fig. 7 in §A.3.1.
Figure 3. Validation accuracy for FL and WW-FL for 500 iterations (top: LeNet/MNIST, bottom: ResNet9/CIFAR10).
Q2 – Accuracy Impact of MPC. In Fig. 4, we compare the plaintext validation accuracy (cf. Q1) to our CrypTen simulation to measure the impact of MPC (i.e., fixed-point representation with 22 bit decimal representation and truncation). While there is a slight difference in initial rounds, both implementations quickly converge to almost the same validation accuracy, with only a small difference on the order of 0.1%.
Figure 4. Validation accuracy for FL and WW-FL in plaintext and MPC (CrypTen simulation) for LeNet/MNIST training.
Q3 – MPC Overhead. Finally, we evaluate the overhead of MPC for secure training and aggregation. For this, we measure the run-times and communication for one iteration of LeNet/MNIST training (i.e., 5 local epochs) in AWS for one cluster (with 1Gbps bandwidth and 20ms latency) and one iteration of global aggregation (with 100Mbps bandwidth and 100ms latency). The training on the cluster level takes 315s and requires 5.28GB inter-server communication, which is multiple orders of magnitude more than local plaintext training in PyTorch (which only takes 0.07s). The aggregation over 10 cluster inputs is very efficient with 0.023s run-time and has no communication overhead since only linear operations are required, which are local operations in MPC.
@@
5. Attacks & Mitigations
@@
In FL, malicious participants can degrade accuracy through data or model poisoning attacks (Tian et al., 2022). Because WW-FL models are not accessible to clients, these attackers are limited to manipulating the training data they provide, resulting in either targeted, backdoor, or untargeted attacks. Our work is focused on untargeted FL poisoning, as it is most relevant to real-world applications and more difficult to detect (Shejwalkar et al., 2022; Fang et al., 2020). We therefore propose how to systematically evaluate the effectiveness of state-of-the-art data-poisoning attacks (Xiao et al., 2012; Fang et al., 2020; Shejwalkar et al., 2022; Tolpegin et al., 2020) in the WW-FL setting and propose a new robust aggregation scheme as a possible mitigation.
5.1. Data-Poisoning Attacks in WW-FL
@@
In data-poisoning attacks, malicious clients can perform arbitrary manipulations on the training data. State-of-the-art attacks are based on label flipping, where clients keep the legitimate training samples, yet exchange the associated labels according to different strategies (Fang et al., 2020; Tolpegin et al., 2020; Shejwalkar et al., 2022).
@@
[…]
@@
    Targeted Label Flipping (TLF (Tolpegin et al., 2020)): Simply flips all labels from a source class to a target class. In WW-FL, we always set the source class as 0 and the target class as 1.
@@
    Static Label Flipping (SLF (Fang et al., 2020; Shejwalkar et al., 2022)): Uses a fixed permutation to determine the new label for each poisoned sample, given by the equation $\textit{new}\_\textit{label}=\textit{num}\_\textit{classes}-\textit{old}\_\textit{label}-1$.
@@
    Dynamic Label Flipping (DLF (Shejwalkar et al., 2022)): Utilizing a surrogate model to flip labels for each sample, our implementation aggregates data from all malicious clients to train a model with the same architecture as the one used in WW-FL training. Once trained, this model is utilized for inference and the labels are assigned to the least probable output determined by the surrogate model. The term “dynamic” is used as the labels rely on the trained model, and by altering the training configuration, the poisoned labels will be modified accordingly. The exact training settings of the surrogate model are given in Tab. 7.
    Table 7. Training parameters for surrogate model in DLF.
@@
5.2. New Robust Aggregation Scheme
The most common FL aggregation scheme, FedAvg, simply computes a (weighted) average of all inputs (cf. §2). In contrast, robust aggregation schemes detect and exclude outliers, thus a suitable mitigation against data poisoning. From the schemes surveyed in (Shejwalkar et al., 2022; Gehlhar et al., 2023), we identify “FLTrust” (Cao et al., 2021) and “Trimmed Mean” (TM) (Yin et al., 2018) as the ones with most efficient MPC implementations.
FLTrust (Cao et al., 2021) measures the cosine similarity between the inputs of participants and the most recent global model trained with a clean data set, then excludes the least similar inputs. Trimmed Mean (TM) (Yin et al., 2018), for each coordinate, computes the mean across the provided gradient updates and excludes the values that deviate the most in either direction of the mean. Performing TM aggregation obliviously in MPC requires implementing costly sorting to determine the ranking in each coordinate.
We observe that, intuitively, data poisoning in contrast to model poisoning does not result in specific coordinates producing extreme outliers. Hence, we propose a heuristic “Trimmed Mean Variant” that computes the mean and ranking only for a small randomly sampled subset of coordinates. Then, during aggregation, it excludes those inputs that occurred the most as outliers in the sample. We detail our algorithm in Alg. 3, where the underlying MPC functionalities TM-List and TopK-Hitter are as follows:
@@ 5.2. New Robust Aggregation Scheme
0:  $\mathcal{W}=\{W^{i}\}_{i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]}$, ${\alpha}$, ${\beta}$, ${\gamma}$ ${\gamma}=|W^{i}|$, ${\alpha}$ – trim threshold, ${\beta}$ – sample size
0:  ${\langle W_{AGG}\rangle}$ aggregated model after removing outliers
1:  initialize: $\mathcal{Z}\leftarrow\emptyset$ set of outliers
2:  $\mathcal{I}\leftarrow$ sample random ${\beta}$ indices from $[1,{\gamma}]$.
3:  $\mathcal{U}\leftarrow\textsc{TM-List}(\mathcal{W}^{\mathcal{I}},{\alpha})$ $\mathcal{W}^{\mathcal{I}}$–Truncated $\mathcal{W}$ with only indices in $\mathcal{I}$, TM-List performs Trimmed Mean algorithm and returns $2{\alpha}$ outlier values (top and bottom ${\alpha}$) for each index in $\mathcal{I}$, $|\mathcal{U}|=2{\alpha}{\beta}$
4:  $\mathcal{V}\leftarrow\textsc{TopK-Hitter}(\mathcal{U},2{\alpha})$ returns list of $2{\alpha}$ indices that occur most frequently in $\mathcal{U}$
@@
8:  ${\langle W_{AGG}\rangle}\leftarrow\textsc{Agg}(\mathcal{W}\setminus\mathcal{Z})$
@@
Implementation Details
CrypTen lacks a built-in oblivious sorting functionality, so we implement privacy-preserving sorting for so-called CrypTensors. Sorting is necessary to compute trimmed mean and our optimized trimmed mean variant. We minimize the number of comparisons by implementing a bitonic sorting network that generates the necessary comparisons between elements to allow for a non-stable sorting algorithm. For trimmed mean, it is not necessary to preserve the relative order of same valued keys as each of them would have been seen as suspicious anyway. The comparisons are stored in plaintext, as they do not reveal any information about the data. The comparison steps are only computed once and then executed in parallel for each coordinate. For 100 elements, we perform 1077 comparisons and for 10 elements only 31. As the result of each comparison is hidden, we perform the swap operation for each pairs as described in Lst. LABEL:listing:crypten-sorting.

### v4

@@ 4.1 Implementation
We implement WW-FL based on the CrypTen MPC framework developed by Meta [KVH+21]. CrypTen provides a PyTorch-style interface and implements MPC operations with GPU support. Specifically, it implements semi-honest arithmetic and Boolean secure two- and multi-party computation protocols that use a trusted third “helper” party to generate correlated randomness. CrypTen provides a “simulation” mode where the specified computation is performed on a single node in plaintext yet simulates all effects that computation in MPC would have on the results (e.g., due to limited fixed-point precision and truncation). We leverage this mode to efficiently evaluate WW-FL’s accuracy and later the impact of data-poisoning attacks. For our run-time and communication measurements, we benchmark the full MPC protocols. In all our experiments, the fixed-point precision in CrypTen is set to 22 decimal bits (the maximum value recommended by the developers).
@@
Moving our prototype implementation to production requires hardening both the cryptographic and system layers. For example, all processing of sensitive data (even when secret-shared) must happen in constant time to protect against timing-based side-channel attacks. Network connections must be secured, e.g., via mutually authenticated TLS to instantiate communication channels that are assumed to be perfectly secure. Finally, servers must be hardened, their APIs must be secured, and reliable key and identity management must be implemented. Please refer to recommended practice for MPC-based systems defined by ITU-T X.1770 [ITU21] and IEEE P2842 [IEE21] for details.
Setup
Plaintext FL and CrypTen-based WW-FL simulations are run on a single computing platform with two Intel Xeon Platinum 8168 CPUs, 1.5TB RAM, and 16 NVIDIA Tesla V100 GPUs. For realistic results in a distributed MPC deployment with two computational and one helper party, we use three Amazon AWS g3s.xlarge instances with 4 vCPUs, and 8GB GPU memory on an NVIDIA Tesla M60. These instances are located in the same AWS availability zone (due to high costs associated with routing traffic between different zones), yet we simulate intra- and inter-continental network connections by setting the bandwidth to 1Gbps/100Mbps and latency to 20ms/100ms respectively.
Tasks
Following prior work such as [RSWP23, BMP+24, TXLZ25], we evaluate WW-FL on two standard image classification tasks: recognizing (a) hand-written digits using LeNet [AAG+25] trained on MNIST [LBBH98] dataset, and (b) objects in one of 10 classes using ResNet9 trained on CIFAR10 [Kri09] dataset. We simulate 1000 clients overall from which 100 are randomly sampled during each training iteration. In the FL setting, all sampled clients provide their computed update directly to the aggregation server. In the WW-FL setting, each of the 10 MPC clusters has 100 associated clients from which each samples 10 clients at random. Each client has 200 data points assigned to it at random with duplicates allowed between clients.
For plain FL, we use batch size 8 and learning rate 0.005, and train locally for 5 epochs before central aggregation. To allow for a fair comparison between FL and WW-FL, we use a correspondingly scaled batch size of 80 and a learning rate of 0.05 as in [GDG+17]. Each client and MPC cluster performs 5 local training epochs with either a batch size of 8 or 80, respectively, following the recommendation in [GDG+17]. Tab. 7 summarises all the (hyper) parameters used during the training.
Table 7: Training parameters used in FL and $W$-FL.
Parameter FL $W$-FL
@@ 4.2 Results
    What is the accuracy difference between FL [MMR+17] and WW-FL (in plaintext)?
@@
    What is the impact on accuracy for WW-FL when moving from plaintext to MPC?
@@
  • Q4:
    How does WW-FL scale and compare to existing fully private training approaches?
Q1 – FL vs WW-FL
In Fig. 3, we compare the validation accuracy of FL and WW-FL for image classification tasks for 500 and 2000 rounds. Here, WW-FL converges significantly faster than regular FL, e.g., after 500 rounds of training ResNet9 on CIFAR10, WW-FL reaches 85.68% validation accuracy, whereas regular FL only reaches 65.95%. Similarly, for MNIST, WW-FL achieves 98.95% accuracy compared to 98.72% for regular FL. We attribute this to WW-FL pooling training data at the cluster level and thus being able to exploit the known benefits of batching [GDG+17, BCN18, WWCJ22].
(a) LeNet trained on MNIST for 500 epochs.
(b) ResNet9 on CIFAR10 for 500 epochs.
(c) LeNet trained on MNIST for 2000 epochs.
(d) ResNet9 on CIFAR10 for 2000 epochs.
Figure 3: Validation accuracy for FL and $W$-FL.
Q2 – Accuracy Impact of MPC
In Fig. 4, we compare the plaintext validation accuracy (cf. Q1) to our CrypTen simulation to measure the impact of MPC (i.e., fixed-point representation with 22 bit decimal representation and truncation). While there is a slight difference in initial rounds, both implementations quickly converge to almost the same validation accuracy, with only a small difference on the order of 0.1%.
Figure 4: Validation accuracy for FL and WW-FL in plaintext and MPC (CrypTen simulation) for LeNet/MNIST training.
Q3 – MPC Overhead
Finally, we evaluate the overhead of MPC for secure training and aggregation. For this, we measure the run-times and communication for one iteration of LeNet/MNIST training (i.e., 5 local epochs) in AWS for one cluster (with 1Gbps bandwidth and 20ms latency) and one iteration of global aggregation (with 100Mbps bandwidth and 100ms latency). The training for WW-FL on the cluster level takes 315s and requires 5.28GB inter-server communication, which is multiple orders of magnitude more than local plaintext training in PyTorch (which only takes 0.07s). The aggregation over 10 cluster inputs is very efficient with 0.023s run-time and has no communication overhead since only linear operations are required, which are local operations in MPC.
@@
Q4 – Scalability & Comparison
The only existing solutions for privacy-preserving training and inference with cryptographic guarantees and global model privacy are based on either MPC or MHE (cf. §2). Existing MPC-based PPML solutions with global model privacy can be seen as a special case of WW-FL with a single MPC cluster on Layer II and no aggregation at Layer I (cf. Tab. 5). In terms of MHE, Spindle [FTP+21], Poseidon [SPT+21], and Hercules [XHX+23] are currently the most promising solutions for collaborative training and inference with global model privacy. We therefore compare scalability and especially training overheads with these existing approaches.
[…]
For the comparison with MHE, we refer to the experiments reported in [FTP+21, SPT+21, XHX+23]. Spindle’s [FTP+21] communication overhead is demonstrated to scale linearly in the number of data providers (Layer III clients in our terminology) and ciphertext size, which in turn depends on the number and dimension of training samples. This is similar to WW-FL, however, we can easily handle additional load from new clients by adding MPC training clusters, thus not increasing the load for existing clusters or clients. Furthermore, Spindle only supports linear and logistic regression models, whereas we validate WW-FL with more complex convolutional neural networks.
Poseidon [SPT+21] extends and improves upon Spindle: while it also scales linearly in the number of clients and training samples, it supports various neural network architectures. However, it relies on approximations for non-linear layers due to the restrictions of HE schemes. Specifically, the evaluation includes a 3-layer fully-connected neural network for MNIST classification that achieves 89.9% accuracy after 1000 epochs, whereas we validate WW-FL with a network that includes three convolutional and two fully-connected layers (along with non-linear activation functions and pooling layers) [AAG+25], and surpasses 90% accuracy in less than 100 rounds with 5 local epochs in each round (cf. results reported for Q1). More comparable is the evaluation of [SPT+21] on a model for CIFAR10 classification with two convolution and two fully-connected layers, which we therefore reference in our comparison in Tab. 8. Another notable benchmark of Poseidon [SPT+21] is training the three-layer SecureML neural network [MZ17] with three clients for 15 epochs on MNIST, which takes 73.1 hours, whereas MPC frameworks such as Falcon [WTB+21] only require 0.56 hours for the same task.
Hercules [XHX+23] advances the state of the art over Poseidon by enabling more efficient homomorphic matrix operations together with polynomial approximations for non-linear activation functions. It further demonstrates the feasibility of training with up to 50 clients, achieving secure aggregation without relying on multiple non-colluding servers and providing strong privacy guarantees under the passive security model. These results highlight the promise of multi-party homomorphic encryption (MHE)-based approaches for privacy-preserving federated learning. Nevertheless, MHE-based solutions also bring certain costs: they require a multi-party key setup among all clients, entail sizable ciphertexts that increase storage requirements, and typically only support passive security while lacking dedicated mechanisms against poisoning attacks. For instance, training with 10 clients in Hercules on MNIST requires a total storage of 49.06GB across the clients. By contrast, WW-FL adopts an orthogonal design point: clients only store their training data in plaintext, while servers hold additive shares of training data and model parameters, resulting in less than 4GB of storage in total. On the downside, however, WW-FL using naive MPC alone requires the availability of at least two non-colluding servers in every MPC cluster. Moreover, while private training is efficiently parallelized across clusters, the overhead compared to cleartext federated learning becomes significant when considering the overall cost of achieving global model privacy.
In Tab. 8, we summarize performance results for the approaches compared above. For the total run-time and communication overhead for WW-FL, we count data sharing between Layer III clients and Layer II cluster servers, Layer II training, secret-shared parameter up- and download between Layer II and Layer I cluster servers, and secure aggregation on Layer I. For the centralized MPC-based method, we only have to consider data sharing between clients and one MPC cluster with the same configuration as a WW-FL Layer II cluster that executes all training epochs on its own. While this results in slightly reduced communication overhead compared to WW-FL (as secure aggregation can be omitted), there is an order of magnitude increase in run-time as the training procedure cannot be parallelized as for WW-FL. Also, note that it might not be practically feasible for a single cluster to handle the substantial total communication overhead that is evenly distributed among all clusters in WW-FL.
[…]
Table 8: Comparison of total run-time in hours, communication in terabytes, and storage in gigabytes for WW-FL and central MPC-based training for 1000 clients, and MHE-based approaches (Poseidon [SPT+21] and Hercules [XHX+23]) for 10 and 50 clients. Note that the neural networks trained for Poseidon and Hercules are less complex and do not include non-linear layers. Type of layers: CV – convolution, MP – max pooling, FC – fully connected, AP – average pooling.
Metric $W$-FL MPC [KVH+21] [SPT+21] [XHX+23] [SPT+21] [XHX+23]
# clients 1,000 1,000 10 10 50 50
# clients per round 100 100 10 10 50 50
# epochs in total 2,500 2,500 1,000 1,000 25,000 25,000
Accuracy 98.95% 98.95% 88.7% 91.8% 51.8% 54.3%
Model LeNet [AAG+25]: “MNIST” [XHX+23]: “CIFAR-10-N1” [XHX+23]:
3CV, 4ReLU, 2MP, 2FC 3FC 2CV, 1AP, 1MP, 2FC
Total run-time (in h) 43.83 437.57 1.24 0.43 126.26 40.73
Total comm. (in TB) 26.45 26.40 7.03 0.18 3,076.17 102.54
Total storage (in GB) 3.39 3.33 3,140.25 49.06 105,468.5 823.50
5 Attacks & Mitigations
@@
In FL, malicious participants can degrade accuracy through data or model poisoning attacks [TCLY22]. Because WW-FL models are not accessible to clients, these attackers are limited to manipulating the training data they provide, resulting in either targeted, backdoor, or untargeted attacks. Our work is focused on untargeted FL poisoning, as it is most relevant to real-world applications and more difficult to detect [FCJG20, SHKR22, RSWP23, TXLZ25]. We therefore propose how to systematically evaluate the effectiveness of state-of-the-art data-poisoning attacks [XXE12, FCJG20, TTGL20, SHKR22] in the WW-FL setting and propose a new robust aggregation scheme as a possible mitigation.
5.1 Data-Poisoning Attacks in WW-FL
@@
In data-poisoning attacks, malicious clients can perform arbitrary manipulations on the training data. State-of-the-art attacks are based on label flipping, where clients keep the legitimate training samples, yet exchange the associated labels according to different strategies [FCJG20, TTGL20, SHKR22].
@@
    Random Label Flipping (RLF [XXE12]): Each poisoned sample is assigned a random class label, i.e., for $\textit{num}\_\textit{classes}$ classes in the dataset, define the new label as $\textit{new}\_\textit{label}=\texttt{randint}(0,\textit{num}\_\textit{classes}-1)$.
@@
    Targeted Label Flipping (TLF [TTGL20]): Simply flips all labels from a source class to a target class. In WW-FL, we always set the source class as 0 and the target class as 1.
@@
    Static Label Flipping (SLF [FCJG20, SHKR22]): Uses a fixed permutation to determine the new label for each poisoned sample, given by the equation $\textit{new}\_\textit{label}=\textit{num}\_\textit{classes}-\textit{old}\_\textit{label}-1$.
@@
    Dynamic Label Flipping (DLF [SHKR22]): Utilizing a surrogate model to flip labels for each sample, our implementation aggregates data from all malicious clients to train a model with the same architecture as the one used in WW-FL training. Once trained, this model is utilized for inference and the labels are assigned to the least probable output determined by the surrogate model. The term “dynamic” is used as the labels rely on the trained model, and by altering the training configuration, the poisoned labels will be modified accordingly. The exact training settings of the surrogate model are given in Tab. 9.
    Table 9: Training parameters for surrogate model in DLF.
@@
5.2 New Robust Aggregation Scheme
The most common FL aggregation scheme, “FedAvg”, simply computes a (weighted) average of all inputs (cf. §2). In contrast, robust aggregation schemes detect and exclude outliers, thus a suitable mitigation against data poisoning. From the schemes surveyed in [SHKR22, GMS+23, XFG25], we identify “FLTrust” [CFLG21] and “Trimmed Mean” (TM) [YCRB18] as the ones with most efficient MPC implementations.
FLTrust [CFLG21] measures the cosine similarity between the inputs of participants and the most recent global model trained with a clean data set, then excludes the least similar inputs. Trimmed Mean (TM) [YCRB18], for each coordinate, computes the mean across the provided gradient updates and excludes the values that deviate the most in either direction of the mean. Performing TM aggregation obliviously in MPC requires implementing costly sorting to determine the ranking in each coordinate.
We observe that, intuitively, data poisoning in contrast to model poisoning does not result in specific coordinates producing extreme outliers. Hence, we propose a heuristic “Trimmed Mean Variant” that computes the mean and ranking only for a small randomly sampled subset of coordinates. Then, during aggregation, it excludes those inputs that occurred the most as outliers in the sample. We detail our algorithm in Alg. 4, where the underlying MPC functionalities TM-List and TopK-Hitter (cf. Tab. 3) are as follows:
@@ 5.2 New Robust Aggregation Scheme
0:  $\mathcal{W}=\{W^{i}\}_{i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}]}$, $\alpha$}, $\beta$}, $\gamma$} ${\gamma}=|W^{i}|$, $\alpha$} – trim threshold, $\beta$} – sample size
0:  $\langle W_{AGG}\rangle$} # aggregated model after removing outliers
1:  initialize: $\mathcal{Z}\leftarrow\emptyset$ # set of outliers
2:  $\mathcal{I}\leftarrow$ sample random $\beta$} indices from $1,{\gamma}]$.
3:  $\mathcal{U}\leftarrow\textnormal{{TM-List}}(\mathcal{W}^{\mathcal{I}},{\alpha})$ $\mathcal{W}^{\mathcal{I}}$–Truncated $\mathcal{W}$ with only indices in $\mathcal{I}$, TM-List performs Trimmed Mean algorithm and returns $2{\alpha}$ outlier values (top and bottom $\alpha$}) for each index in $\mathcal{I}$, $|\mathcal{U}|=2{\alpha}{\beta}$
4:  $\mathcal{V}\leftarrow\textnormal{{TopK-Hitter}}(\mathcal{U},2{\alpha})$ # returns list of $2{\alpha}$ indices that occur most frequently in $\mathcal{U}$
@@
8:  ${\langle W_{AGG}\rangle}\leftarrow\textnormal{{Agg}}(\mathcal{W}\setminus\mathcal{Z})$
@@
5.3 CrypTen Implementation Details for Trimmed Mean
CrypTen lacks a built-in oblivious sorting functionality, so we implement privacy-preserving sorting for so-called CrypTensors. Sorting is necessary to compute trimmed mean and our optimized trimmed mean variant. We minimize the number of comparisons by implementing a bitonic sorting network that generates the necessary comparisons between elements to allow for a non-stable sorting algorithm. For trimmed mean, it is not necessary to preserve the relative order of same valued keys as each of them would have been seen as suspicious anyway. The comparisons are stored in plaintext, as they do not reveal any information about the data. The comparison steps are only computed once and then executed in parallel for each coordinate. For 100 elements, we perform 1077 comparisons and for 10 elements only 31. As the result of each comparison is hidden, we perform the swap operation for each pairs as described in Lst. 1.